it takes 7 painters to paint 3 homes . how many painterss can paint 10 homes
do I cross multiply ?

Answer by stanbon(75887)

it takes 7 painters to paint 3 homes . how many painters can paint 10 homes
---
Yes, after you set up the proportion:
painters/homes = painters/homes
----
p/10 = 7/3
---
p = 10(7/3)
---
p = 70/3 = 23 1/3 painters
=============================
Cheers,
Stan H.
==============